The sum of three numbers is 138. The second is 5 more than the smallest and the third is 10 more than the smallest. Find the numbers. Thank you
Answer by mananth(16946)

The sum of three numbers is 138. The second is 5 more than the smallest and the third is 10 more than the smallest. Find the numbers. Thank you
Let the smallest number be x
second will be x+5
third will be x+10
x+x+5+x+10=138
3x+15=138
3x=138-15
3x=123
x=123/3
=41
the numbers are 41, 46 and 51